OpenStudy (anonymous):

The sides of the cubes are numbered 1 through 6. If they are both tossed, what is the probability that they will both be a 2?

OpenStudy (saifoo.khan):

o.O

OpenStudy (amistre64):

2/36 = 1/18 maybe?

OpenStudy (llort):

(1/6)^2=1/36

OpenStudy (amistre64):

2,2 2,2 are both in the sample space; 2 out of 36 = 1 out of 18

OpenStudy (llort):

but they are not dependent on one another... it is (1/6)*(1/6)

OpenStudy (amistre64):

the question is worded poorly to begin with :)

OpenStudy (amistre64):

are there 2 dice? or one dices?

OpenStudy (anonymous):

2

OpenStudy (amistre64):

and they are tossed once and you wanna know if you get a 2 and a 2 correct?

OpenStudy (anonymous):

you were correct with 1/36 that was one of the answer choices

OpenStudy (amistre64):

being one fo the answer choices doesnt imply correctness :) if you picked it and it said you were correct, that will do lol